OpenStudy (anonymous):

How do I find the value of Sigma n=1 to infinity n/k^n (k>1)

OpenStudy (anonymous):

n = 1 to infinity or k = 1 to infinity?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} n/k ^{2}\] n=1 to infinity... k>1

OpenStudy (anonymous):

i am confused because for fixed k this sum does not converge. for example if k = 1 you get \[\Sigma_1^\infty n\] which is infinite

OpenStudy (anonymous):

It's supposed to be solved using the results of the 1st question I asked involving showing the \[\sum_{1}^{\infty} n x ^{n} = f(x) x/(1-x)^2\]